what is the area of a rectangle with perimeter 120 ft and a width that is half as long as its length
Answer by ewatrrr(24785)

 
Hi,

P = 2L +2w + 120ft
L + w = 60ft
L + L/2 = 60ft
2L + L = 120
3L = 120
L = 40ft and w = 20ft